# A Learning Framework for Diverse Legged Robot Locomotion Using Barrier-Based Style Rewards

### v3

@@ Abstract
footnotetext: This research was supported in part by the Challengeable Future Defense Technology Research and Development Program through the Agency For Defense Development(ADD) funded by the Defense Acquisition Program Administration(DAPA) in 2024(No.912768601), and in part by the Technology Innovation Program(or Industrial Strategic Technology Development Program-Robot Industry Technology Development)(00427719, Dexterous and Agile Humanoid Robots for Industrial Applications) funded by the Ministry of Trade Industry & Energy(MOTIE, Korea).11footnotetext: Gijeong Kim, Yong-Hoon Lee, and Hae-Won Park are with the Department of Mechanical Engineering, Korea Advanced Institute of Science and Technology, Yuseong-gu, Daejeon 34141, Republic of Korea. haewonpark@kaist.ac.kr**footnotetext: These authors contributed equally to this work\$$\dagger$\$$\dagger$\$footnotetext: Corresponding authorfootnotetext: Supplementary video: https://youtu.be/JV2_HfTlOKI
@@ I INTRODUCTION
To address the challenge of complex reward tuning in RL, several approaches have been explored to produce natural locomotion in legged robots. By leveraging prior motion data from animal motion capture or trajectory optimization, natural gait motions in RL can be achieved through the addition of imitation reward terms [9, 10, 11] or imitation frameworks [12]. Alternatively, without prior motion data, using gait-specific rewards can enhance the versatility of locomotion (walking, running, hopping, skipping) in biped robots with a predefined preferred gait [13]. Recently, Kim et al. [14] demonstrated that applying constraints in constrained RL to define desirable operational motion ranges reduces reward engineering efforts and enhances the framework’s generalizability across different robot types.
Here, we introduce a learning framework that explicitly guides motion style with a preferred gait, rather than strictly constraining specific motion characteristics or gait sequences, to enhance task versatility for legged robots while facilitating animal-like natural motion. To achieve this flexible guidance, we propose a novel reward function based on the relaxed-logarithmic barrier function, a soft-constraint type originally used in trajectory optimization [15]. Unlike constrained RL approaches that prioritize safe learning and strict constraint satisfaction [16], our method focuses on guiding motion characteristics while maintaining flexibility. The preferred gait is also shaped by our proposed reward function, which features an explicitly tunable margin that facilitates adjustments of stance and swing times based on task requirements.
@@ II Related Works
Guiding to Desired Motion Style  To facilitate natural motion without complex reward engineering, previous studies have incorporated additional reward terms to define motion styles: if a prior dataset is available, discriminator-based style rewards [18, 9, 19, 10] or direct reward formulations (e.g., reducing discrepancies between reference and resultant motion) [20, 11] can be used to imitate prior motions. However, this approach can be limited for hard-to-obtain data, such as tripod locomotion or biped locomotion of quadruped animals. In an alternative approach, Kim et al. [14] explicitly define locomotion styles by imposing constraints on physical quantities, such as joint angles and velocities, in constrained RL. They employ Interior Point Optimization (IPO)[21], which manages constraints with logarithmic barrier functions. This requires additional algorithms to handle the infinity regions during constraint violations[14, 16]. Our approach also explicitly defines locomotion styles through constraints; however, unlike previous methods, we utilize soft constraints within the reward functions, rather than cost functions in constrained RL. By employing a relaxed logarithmic barrier [15], we eliminate the need for additional algorithms to manage infinity regions of the log-function, simplifying the implementation.
@@ III-B Gait Encoding and Foot Clearance
Task-dependent Gait  Task-specific gaits are set by adjusting the gait period $T$ and phase offsets $\phi_{i}$ for each leg, denoted as RH, LH, RF, LF for $i=1,2,3,4$, respectively. For the quadruped mode traversing rough terrain, we employ a trotting gait with $T=0.72\,\text{sec}$, setting diagonal leg pairs $\phi_{1,3}=0$ and $\phi_{2,4}=0.5\,T$. Based on this basic trot setting, tripod and biped modes are implemented: 1) In tripod mode, only the constraint variables $f_{i^{\prime}}$ (Eq. 2) and $l_{i^{\prime}}$ (Eq. 3) are adjusted for the non-contacting $i^{\prime}$-th foot. Here, $f_{i^{\prime}}$ is set to $-1$ if contacting and $1$ if not (for non-contact), while $l_{i^{\prime}}$ is set to always adhere to the first case in Eq. 3 (for continuous foot lift). 2) In biped mode, constraints for gait and foot clearance are enforced only for the two contacting feet.
In quadruped mode for overcoming high steps, we employ a transverse gallop gait with $T=0.82\,\text{sec}$ and phase offsets $[0,0.2\,T,0.5\,T,0.7\,T]$ for $\phi_{1},\phi_{2},\phi_{3},\phi_{4}$ respectively. For fast running tasks, we use a rotary gallop gait with $T=0.72\,\text{sec}$ and phase offsets $[0,0.2\,T,0.7\,T,0.5\,T]$. Using the heuristics from [30], for command velocities above 1.75 m/s, stance time $T_{\text{st}}$ is computed as $L/v^{\text{cmd}}_{x}$ where $L$ is the stride length (0.63 m) and $v^{\text{cmd}}_{x}$ is the forward velocity command. Swing time $T_{\text{st}}$ remains fixed at 0.36 sec across all velocities. Phase offsets are set as $[0,0.4\,T_{\text{st}},0.5\,T+0.4\,T_{\text{st}},0.5\,T]$, ensuring continuous adaptation at 1.75 m/s.
@@ III-C Mode-dependent Barrier Reward Design
The parameter $\delta$ determines the severity of the penalty in the constraint-violation region; smaller $\delta$ values yield steeper gradients. This setting should align with each constraint’s operational range, as identical values of deviation can have different impacts across physical quantities. For instance, a 1 rad deviation in joint position can be considered more severe than a 1 rad/s deviation in joint velocity. Additionally, $\delta$ balances the enforcement among different constraint rewards based on their severity, functioning implicitly as a weight.
The fine tuning is flexible for each task. For instance, in high step climbing, body height constraints are widened to allow a steeper body pitch for climbing maneuvers; in fast running with galloping, the lower boundary for gait is set to $-0.3$ for stricter gait enforcement, joint velocity boundaries are adjusted to ±$\pm 16$ rad/s, and pitch velocity boundaries are expanded for command speed exceeding $3.0$ m/s.
@@ IV-A Comparison with the Baseline
Comparison 1: Learning to Walk  We evaluated the ability to track forward velocity commands without falling in the simulation as learning iterations progressed. The robot was initially oriented in the +x direction, and the command increased from 0 to 1 m/s during the initial 1 second and was maintained for the following 9 seconds. This resulted in a target position of $+9.5$ m along the x-axis, as indicated by red lines in Fig. 4. Fig. 4(a) and (b) represent the robot’s final position along the x-axis in quadruped and biped modes, respectively. Our framework learns to maintain balance and walk earlier than the baseline, reaching the target in fewer iterations (within 1,000 iterations), whereas the baseline could not reach it within the given range of 2,500 iterations. The efficacy of the initial learning stages is attributed to the barrier reward. The initialized policy frequently explores the constraint-violation region and, facilitated by steep gradients, efficiently learns the desired motion style and preferred gait, while also being guided to follow the command velocity.
Comparison 2: Climbing High Step  For the high-step task, curriculum learning gradually increases step heights to 69 cm over 7,600 iterations in a consecutive-step environment. Fig. 4(c) compares the step-climbing performance of our framework to that of the baseline. Max-Height is defined as the highest step height at which the success rate of climbing three steps within the time limit exceeds 50 % in 100 randomized environments. At 3,000 iterations, our framework consistently overcame the highest obstacles in the training environment for commands ranging from 0.3 m/s to 1.5 m/s. In contrast, the baseline failed to overcome obstacles at low-speed commands. As iterations progressed, our framework increased Max-Height across most commands, aided by our explicit enforcement of foot clearance within the barrier reward. In contrast, the baseline exhibited performance degradation due to unstable foot clearance patterns, beginning with low speeds and ultimately leading to a non-moving policy.
Figure 4: We compare the performance of our method against the baseline in (a) Quadruped locomotion on flat terrain (b) Biped locomotion on flat terrain (c) Quadruped locomotion for overcoming high steps. 10 policies trained with different random seeds, were evaluated for both our method and the baseline across 100 random initial conditions for each policy. Each graph shows the mean and standard deviation.
@@ IV-B Real-World Experiments
Climbing High Step  Our method demonstrated the ability to overcome 58 cm and 67 cm obstacles in hardware experiments using HOUND and HOUND2, respectively. To our knowledge, this achievement represents the highest obstacle clearance in blind locomotion of a quadruped robot. Considering larger robots’ advantage in overcoming higher steps, we defined a Step-to-Leg ratio by dividing the maximum step height by the combined leg lengths (thigh and calf) for fair comparisons. Lee et al. [8], Nahrendra et al. [32], and Wu et al. [10] achieved step heights of 16.8 cm (Step-to-Leg 0.27) with the ANYmal, 15 cm (Step-to-Leg 0.38) with the Unitree A1, and 25 cm (Step-to-Leg 0.59) with the Unitree Go1, respectively. Our results for both absolute step height (67 cm) and the Step-to-Leg ratio (0.96 for HOUND2) outperform those of existing RL-based blind locomotion studies on quadruped robots.
@@
Biped Mode  We demonstrated that the quadruped robot HOUND can walk and run in biped mode, following commanded velocities. Even with point feet, the robot resists hard pushes from both front and back by using the pitch motion of the hip joint. Notably, it also withstands lateral pushes, even without an actuated DOF for roll motion in its biped configuration. By using the same heuristics of varying $T_{\text{st}}$ from galloping tasks, we achieved a bipedal locomotion speed of 3.6 m/s (Froude Number = 1.40) based on treadmill measurements. This is faster than the Unitree G1, which reached 3.3 m/s (Fr = 1.18), and Unitree H1, with a speed of 2.0 m/s (Fr = 0.82). For the stair climbing task, the same stair environment used in quadruped mode is employed during learning. Finally, the robot performs biped locomotion while carrying a box. In the learning stage, the box’s position and mass are added to the privileged state $\mathbf{x}_{t}$, and the robot’s nominal configuration is adjusted to support a box. An additional barrier reward ensures the box is properly positioned above the robot’s arm. In real-world experiments, the robot carries a 7.5 kg load, demonstrating push recovery.
@@ V Conclusion
In this study, we present a learning framework that enables various modes of operation (biped, tripod, quadruped) on a single quadruped robot, utilizing a barrier-based style reward. The framework effectively guides the desired motion style and facilitates adjustments to the preferred gait as required by the task. Extensive real-world experiments across diverse environments demonstrate effectiveness of the framework.

### v4

@@ Abstract
00footnotetext: This research was supported in part by the Challengeable Future Defense Technology Research and Development Program through the Agency For Defense Development(ADD) funded by the Defense Acquisition Program Administration(DAPA) in 2024(No.912768601), and in part by the Technology Innovation Program(or Industrial Strategic Technology Development Program-Robot Industry Technology Development)(00427719, Dexterous and Agile Humanoid Robots for Industrial Applications) funded by the Ministry of Trade Industry & Energy(MOTIE, Korea).00footnotetext: Gijeong Kim, Yong-Hoon Lee, and Hae-Won Park are with the Department of Mechanical Engineering, Korea Advanced Institute of Science and Technology, Yuseong-gu, Daejeon 34141, Republic of Korea. haewonpark@kaist.ac.kr00footnotetext: These authors contributed equally to this work00footnotetext: Corresponding author00footnotetext: Supplementary video: https://youtu.be/JV2_HfTlOKI
@@ I INTRODUCTION
To achieve desirable gait motion while mitigating the challenge of complex reward tuning in RL, several approaches have been explored. By leveraging prior motion data from animal motion capture or trajectory optimization, imitation reward terms [9, 10, 11] or frameworks [12] incorporated in RL can realize targeted motions. Alternatively, without prior motion data, using gait-specific rewards can enhance the versatility of locomotion (walking, running, hopping, skipping) in biped robots with a predefined preferred gait [13]. Recently, Kim et al. [14] demonstrated that applying constraints in constrained RL to define operational motion ranges reduces reward engineering efforts and enhances the framework’s generalizability across different robot types.
Here, we introduce a learning framework that explicitly guides motion style with a preferred gait, rather than strictly constraining specific motion characteristics or gait sequences, to enhance task versatility for legged robots while facilitating animal-like natural motion. To achieve this flexible guidance, we propose to utilize a reward function based on the relaxed-logarithmic barrier function, a soft constraint type originally used in trajectory optimization [15]. Unlike constrained RL approaches that prioritize safe learning and strict constraint satisfaction [16], our method focuses on guiding motion characteristics while maintaining flexibility. The preferred gait is also shaped by our proposed reward function, which features an explicitly tunable margin that facilitates adjustments of stance and swing times based on task requirements.
@@ II Related Works
Guiding to Desired Motion Style  To facilitate natural motion without complex reward engineering, previous studies have incorporated additional reward terms to define motion styles: If a prior dataset is available, discriminator-based style rewards [18, 9, 19, 10] or direct reward formulations (e.g., reducing discrepancies between reference and resultant motion) [20, 11] can be used to imitate prior motions. However, this approach can be limited for hard-to-obtain data, such as tripod locomotion or biped locomotion of quadruped animals. Alternatively, Kim et al. [14] explicitly define locomotion styles by imposing constraints on physical quantities, such as joint angles and velocities, in constrained RL. They employ Interior Point Optimization (IPO)[21], which manages constraints with logarithmic barrier functions. This requires additional algorithms to handle the infinity regions during constraint violations[14, 16]. Our approach explicitly defines locomotion styles through constraints; however, unlike previous methods, we incorporate a relaxed logarithmic barrier [15] into the reward function. This eliminates the need for additional algorithms to handle infinite log-function regions and simplifies implementation.
@@ III-A Relaxed Logarithmic Barrier Reward
In trajectory optimization, Hauser et al. [15] iteratively reduce $\delta$ and $\alpha_{k}$ to ensure convergence to the original constrained optimization problem. In this work, our goal is to guide motion style rather than strictly enforce constraints. Therefore, we keep $\delta$ and $\alpha_{k}$ constant to simplify the training process, as we verified that this setting still effectively achieves the desired motion style.
@@ III-B Gait Encoding and Foot Clearance
Task-dependent Gait  Task-specific gaits are set by adjusting the gait period $T$ and phase offsets $\phi_{i}$ for each leg, denoted as RH, LH, RF, LF for $i=1,2,3,4$, respectively. For the quadruped mode traversing rough terrain, we employ a trotting gait with $T=0.72\,\text{sec}$, setting diagonal leg pairs $\phi_{1,3}=0$ and $\phi_{2,4}=0.5$. Based on this basic trot setting, tripod and biped modes are implemented: 1) In tripod mode, only the constraint variables $f_{i^{\prime}}$ (Eq. 2) and $l_{i^{\prime}}$ (Eq. 3) are adjusted for the non-contacting $i^{\prime}$-th foot. Here, $f_{i^{\prime}}$ is set to $-1$ if contacting and $1$ if not (for non-contact), while $l_{i^{\prime}}$ is set to always adhere to the first case in Eq. 3 (for continuous foot lift). 2) In biped mode, constraints for gait and foot clearance are enforced only for the two contacting feet.
In quadruped mode for overcoming high steps, we employ a transverse gallop gait with $T=0.82\,\text{sec}$ and phase offsets $[0,0.2,0.5,0.7]$ for $\phi_{1},\phi_{2},\phi_{3},\phi_{4}$ respectively. For fast running tasks, we use a rotary gallop gait with $T=0.72\,\text{sec}$ and phase offsets $[0,0.2,0.7,0.5]$. Using the heuristics from [30], for command velocities above 1.75 m/s, stance time $T_{\text{st}}$ is computed as $L/v^{\text{cmd}}_{x}$ where $L$ is the stride length (0.63 m) and $v^{\text{cmd}}_{x}$ is the forward velocity command. Swing time $T_{\text{sw}}$ remains fixed at 0.36 sec across all velocities. Phase offsets are set as $[0,0.4\,\frac{T_{\text{st}}}{T},0.5+0.4\,\frac{T_{\text{st}}}{T},0.5]$, ensuring continuous adaptation at 1.75 m/s.
@@ III-C Mode-dependent Barrier Reward Design
Figure 3: Velocity tracking performance for different $\delta$ values ($1/2$, 1, and 2 times the default setting in Table I) across constraints. Each configuration was evaluated in simulation under randomized conditions identical to the training setup for Traversing Outdoor Terrain in IV-B, where environments ranged from flat ground to challenging steps and slopes. The graph presents the mean tracking error norms of $v_{x}$, $v_{y}$, and $\omega_{z}$ over a 5-second period across 2000 environments with randomized command velocities.
Within these predefined ranges, the parameter $\delta$ determines the severity of penalties in constraint-violation regions; smaller values yield steeper gradients. This setting should align with each constraint’s range; for example, we set $\delta$ to 10–30 % of each constraint’s range for most constraints, using a tighter 4 % for joint position to prevent singular poses and a wider 50 % for base motion to avoid overly sharp penalties that could hinder movement.
The constraints enable task-specific adjustments. For instance, in high step climbing, the body height constraints are widened to allow a steeper body pitch for climbing maneuvers; in fast running with galloping, the lower boundary for gait is set to $-0.3$ for stricter enforcement, joint velocity boundaries are adjusted to ±$\pm 16$ rad/s, and pitch velocity boundaries are expanded for command speed over $3.0$ m/s.
Sensitivity of $\delta$ selection  Fig. 3 presents the velocity tracking performance for different $\delta$ values across constraints. The results indicate that doubling or halving the default $\delta$ in Table I did not lead to noticeable performance differences, suggesting that the proposed method is not too sensitive to $\delta$ selection and can alleviate hyperparameter tuning. However, excessively small or large $\delta$ can cause deviations from the desired motion (e.g., overly strict foot clearance enforcement leading to exaggerated foot lifting).
@@ IV-A Comparison with the Baseline
Comparison 1: Learning to Walk  The training was conducted on flat terrain, and the robot learned to track random velocity commands. During evaluation, the robot aims to accelerate for one second and then maintain a speed of 1 m/s, targeting a position as indicated by the red lines in Fig. 5. Fig. 5(a) and (b) represent the robot’s final position along the x-axis in quadruped and biped modes, respectively. Our framework achieves the target in fewer iterations than the baseline by guiding desirable gait motion and command velocity tracking from the early stages of learning, facilitated by the barrier reward. In contrast, the baseline gradually learns to stand, begins walking, and converges but does not fully reach the target position.
Comparison 2: Climbing High Step  For the high-step task, curriculum learning gradually increases step heights to 69 cm over 7,600 iterations in a consecutive-step environment. Fig. 5(c) compares the step-climbing performance of our framework to that of the baseline. Max-Height is defined as the highest step height at which the success rate of climbing three steps within the time limit exceeds 50 %. At 3,000 iterations, our framework consistently overcame the highest obstacles in the training environment for commands ranging from 0.3 m/s to 1.5 m/s. In contrast, the baseline failed to overcome obstacles at low-speed commands. As iterations progressed, our framework increased Max-Height across most commands, aided by our explicit enforcement of foot clearance within the barrier reward. In contrast, the baseline exhibited performance degradation due to unstable foot clearance patterns, beginning with low speeds and ultimately leading to a stationary policy.
Figure 5: We compare the performance of our method against the baseline in (a) Quadruped locomotion on flat terrain (b) Biped locomotion on flat terrain (c) Quadruped locomotion for overcoming high steps. 10 policies trained with different random seeds, were evaluated for both our method and the baseline across 100 random initial conditions for each policy. Each graph shows the mean and standard deviation. In (a) and (b), both our method and the baseline converge with no further improvement after 2500 iterations.
@@ IV-B Real-World Experiments
Climbing High Step  Our method demonstrated the ability to overcome 58 cm and 67 cm obstacles in hardware experiments using HOUND and HOUND2, respectively. To verify robustness, we conducted five tests for each case, all of which were successful. To our knowledge, this achievement represents the highest obstacle clearance in blind locomotion of a quadruped robot. Considering larger robots’ advantage in overcoming higher steps, we defined a Step-to-Leg ratio by dividing the maximum step height by the combined leg lengths (thigh and calf) for fair comparisons. Lee et al. [8], Nahrendra et al. [32], and Wu et al. [10] achieved step heights of 16.8 cm (Step-to-Leg 0.27) with the ANYmal, 15 cm (Step-to-Leg 0.38) with the Unitree A1, and 25 cm (Step-to-Leg 0.59) with the Unitree Go1, respectively. Our results for both absolute step height (67 cm) and the Step-to-Leg ratio (0.96 for HOUND2) outperform those of existing RL-based blind locomotion studies on quadruped robots.
@@
Biped Mode  We demonstrated that the quadruped robot HOUND can walk and run in biped mode, following commanded velocities. Even with point feet, the robot resists hard pushes from both front and back by using the pitch motion of the hip joint. Notably, it also withstands lateral pushes, even without an actuated DOF for roll motion in its biped configuration. By using the same heuristics of varying $T_{\text{st}}$ from galloping tasks, we achieved a bipedal locomotion speed of 3.6 m/s based on treadmill measurements. For the stair climbing task, the same stair environment used in quadruped mode is employed during learning. Finally, the robot performs biped locomotion while carrying a box. In the learning stage, the box’s position and mass are added to the privileged state $\mathbf{x}_{t}$, and the robot’s nominal configuration is adjusted to support a box. An additional barrier reward ensures the box is properly positioned above the robot’s arm. In real-world experiments, the robot carries a 7.5 kg load, demonstrating push recovery.
@@ V Conclusion
In this study, we present a learning framework that enables various modes of operation (biped, tripod, quadruped) on a single quadruped robot, utilizing a barrier-based style reward. The framework effectively guides the desired motion style and facilitates adjustments to the preferred gait as required by the task. Extensive real-world experiments across diverse environments demonstrate effectiveness of the framework. Future work aims to generalize the barrier formulation to develop a unified policy capable of handling diverse tasks.